### v2

# Probing quantum entanglement with Generalized Parton Distributions at the Electron-Ion Collider

Abstract
Within the collinear factorization framework based on Generalized Parton Distributions (GPDs), we calculate the spin density matrix of exclusively produced quark and antiquark pairs $u\bar{u}$, $d\bar{d}$, $s\bar{s}$, $c\bar{c}$, $b\bar{b}$ in Deep Inelastic Scattering. The presence of both real and imaginary parts in the GPD amplitudes leads to a rich pattern of entanglement between the quark and the antiquark. We map out kinematical regions where the pairs exhibit entanglement, Bell nonlocality and non-stabilizerness (‘magic’). We also predict that massive quarks and antiquarks are transversely polarized, similar to the well-known transverse hyperon polarization in unpolarized collisions. In strangeness, charm and bottom productions, the polarization can reach 50-80% in certain kinematic regions in the low-energy runs of the Electron-Ion Collider.
I Introduction
Entanglement is a quintessential feature of quantum mechanics that requires the coherent description of a composite quantum system as a whole. While it manifests across vastly different scales in both system size and particle number, historically stringent experimental tests of entanglement have primarily focused on the spin or polarization of elementary particle pairs. A classic example is the entanglement of photon polarizations in the (para-)positronium decay ${\rm Ps}\to 2\gamma$ Wu:1950zz. More definitive tests including the violation of Bell’s inequality Bell:1964kc, or the Clauser-Horne-Shimony-Holt (CHSH) inequality also involve two-photon final states Clauser:1969ny; Aspect:1981zz; Weihs:1998gy. Photons are advantageous because their polarizations can be measured relatively easily in table-top or small-scale laboratory experiments.
Entanglement is also ubiquitous among the other elementary particles of the Standard Model that are produced in large-scale collider experiments. However, its direct measurement is much more challenging for a number of reasons. First, many redundant particles are produced in high energy collisions and they may interfere with the particles of interest. Besides, if the strong interaction is involved, entanglement generated at the partonic level can be easily washed out by the hadronization process. Moreover, the polarization of final state particles is not directly measurable. Under these circumstances, in practice, one can only infer entanglement between a pair of heavy, short-lived particles by measuring the angular distribution of their weak decay products. The first experimental confirmation of entanglement in the top-antitop ($t\bar{t}$) system, as predicted by theory Bernreuther:1993hq; Baumgart:2012ay; Bernreuther:2015yna; Afik:2020onf; Dong:2023xiw, has been achieved by the ATLAS and CMS collaborations at the LHC ATLAS:2023fsd; CMS:2024pts; CMS:2024zkc (see also an earlier attempt CMS:2019nrx). Expectations are high Barr:2024djo; Afik:2025ejh that similar measurements can be done for other particle species such as $W^{+}W^{-}$ Aguilar-Saavedra:2022mpg; Fabbrichesi:2023cev, $\tau^{+}\tau^{-}$ Altakach:2022ywa; Han:2025ewp and $b\bar{b}$ Kats:2023zxb; Afik:2025grr pairs, etc., and even light quark-antiquark pairs Cheng:2025cuv. Moreover, a violation of the Bell-CHSH inequality, or ‘Bell nonlocality’ could also be detected in these systems Afik:2025ejh, although it adds another layer of practical challenges.
In this paper, we continue the discussion of spin-spin entanglement between quark-antiquark ($q\bar{q}$) pairs in Deep Inelastic Scattering (DIS) and Ultraperipheral Collisions (UPCs) along the lines of Qi:2025onf; Fucilla:2025kit. Our work is particularly motivated by the future high-luminosity DIS experiment at the Electron-Ion Collider (EIC) AbdulKhalek:2021gbh. Since an electron is involved in the initial state, DIS offers a cleaner (i.e., less produced particles) environment for studying entanglement than in hadron-hadron collisions. One can realize an even cleaner environment, akin to the final states of $e^{+}e^{-}$ annihilation, by focusing on exclusive (diffractive) processes where the target proton does not break up. In Fucilla:2025kit, the spin density matrix of this process has been computed in the Regge limit where the collision energy is asymptotically large and only the color-singlet gluonic exchange (‘Pomeron’) needs to be considered. It has been found that the $q\bar{q}$ pairs are always entangled and always exhibit Bell nonlocality. On the other hand, the EIC runs with variable electron-proton $e+p$ center-of-mass (CM) energies in the range $28<\sqrt{s}<140$ GeV which only partially overlaps with the Regge regime. A solid theoretical framework to describe exclusive processes in this window is the Generalized Parton Distribution (GPD) Diehl:2003ny; Belitsky:2005qn. So far, the study of GPDs has been mostly centered around the goal of obtaining the three dimensional tomographic picture of the nucleon. Here, for the first time, we propose to use GPDs as a tool to probe quantum entanglement at the EIC. We will be particularly interested in whether entanglement and Bell nonlocality are generic features of color singlet exchanges in QCD, or they are emergent phenomena in the high energy limit. We also evaluate the so-called ‘magic’ Leone:2021rzd; White:2024nuc, a necessary nonclassical resource for achieving quantum advantage in quantum computing.
As a byproduct, we find nonvanishing polarization of massive $q\bar{q}$ pairs, even though the colliding particles are unpolarized. This is a kind of single spin asymmetry, and is analogous to the transverse polarization of hyperons in unpolarized proton-proton collisions and DIS Bunce:1976yb; Lundberg:1989hw; HERMES:2007fpi; COMPASS:2021bws. We demonstrate that the polarization can be quite sizable, exceeding 60% in certain kinematical regions.
II Exclusive dijet cross section
In this section, we first review the exclusive dijet cross section calculated in the collinear factorization framework Braun:2005rg. We then derive the associated spin density matrix by explicitly retaining the spin degrees of freedom of the produced quark and antiquark.
To lowest order, ‘dijet’ simply means an energetic quark-antiquark ($q\bar{q}$) pair. ‘Exclusive’ means that that the target proton scatters elastically. Thus the process of interest is $\gamma^{(*)}+p\to q+\bar{q}+p^{\prime}$. This can be thought of as a subprocess in DIS where the virtual photon $\gamma^{*}$ with virtuality $Q^{2}=-q^{2}$ is emitted from the incoming electron. In proton-nucleus ($p+A$) UPCs, the photon $\gamma$ is real ($Q^{2}\approx 0$), emitted from the incoming heavy nucleus. We work in a frame where the photon is collinear [^1] The light-cone coordinates are defined as $p^{\mu}=(p^{+},p^{-},{\bm{p}})$ with $p^{\pm}=\frac{1}{\sqrt{2}}(p^{0}\pm p^{3})$. Boldface letters denote two-dimensional vectors ${\bm{p}}=(p^{1},p^{2})$ and we use $i,j=1,2$ for their indices. $q^{\mu}=(q^{+},-\frac{Q^{2}}{2q^{+}},{\bm{0}})$ and fast-moving $q^{+}\gg Q$ in the $x^{3}$ direction. The incoming and outgoing protons have momenta
$p^{\mu}\approx\left(0,(1+\xi)P^{-},-\frac{{\bm{\Delta}}}{2}\right),\qquad p^{\prime\mu}\approx\left(0,(1-\xi)P^{-},\frac{{\bm{\Delta}}}{2}\right).$ (1)
We will neglect the proton mass and assume ${\bf\Delta}\approx 0$ throughout this paper. The produced $q\bar{q}$ pair has momentum
$\tilde{k}^{\mu}=\left(zq^{+},\frac{k_{\perp}^{2}+m^{2}}{2zq^{+}},{\bm{k}}\right),\qquad\tilde{k}^{\prime\mu}=\left(\bar{z}q^{+},\frac{k_{\perp}^{2}+m^{2}}{2\bar{z}q^{+}},-{\bm{k}}\right),$ (2)
and invariant mass
$M^{2}=(\tilde{k}+\tilde{k}^{\prime})^{2}=\frac{k_{\perp}^{2}+m^{2}}{z\bar{z}}.$ (3)
$0<z=\frac{p\cdot\tilde{k}}{p\cdot q}<1$ is the fraction of the photon energy carried by the quark ($\bar{z}=1-z$ for the antiquark). One may regard this process as the $2\to 2$ scattering $\gamma^{(*)}+{\mathbb{P}}\to q+\bar{q}$ where $\mathbb{P}$} denotes ‘Pomeron’ [^2] The use of the word ‘Pomeron’ is due to a lack of better terminology. In the present context, it simply means the color singlet two-gluon or quark-antiquark exchange in the $t$-channel. In the high energy (Regge) limit, it is smoothly connected to the Pomeron exchange discussed in Fucilla:2025kit. with momentum
${\mathbb{P}}^{\mu}=p^{\mu}-p^{\prime\mu}\approx\left(0,2\xi P^{-},{\bm{0}}\right).$ (4)
The skewness variable $\xi$ is fixed by momentum conservation in this $2\to 2$ scattering as
$\frac{2\xi}{1+\xi}=\frac{M^{2}+Q^{2}}{W^{2}+Q^{2}}=\frac{M^{2}+Q^{2}}{ys},$ (5)
with $W^{2}=(p+q)^{2}$ being the $\gamma^{*}+p$ center-of-mass (CM) energy. On the right hand side, we introduced the $e+p$ CM energy $s$ and the standard DIS variable $y=\frac{q\cdot p}{l\cdot p}$.
We write the cross section in unpolarized DIS as
$\frac{d\sigma^{ep}}{dW^{2}dQ^{2}dzd^{2}{\bm{k}}d^{2}{\bm{\Delta}}}=\frac{\alpha_{em}}{\pi sQ^{2}}\frac{1+(1-y)^{2}}{2y}\left(\frac{d\sigma^{T}}{dzd^{2}\boldsymbol{k}d^{2}{\bm{\Delta}}}+\varepsilon\frac{d\sigma^{L}}{dzd^{2}\boldsymbol{k}d^{2}{\bm{\Delta}}}\right)+\cdots.$ (6)
The superscripts $T$ and $L$ denote the contributions from the transversely and longitudinally polarized virtual photons and $\varepsilon=\frac{1-y}{1-y+\frac{y^{2}}{2}}$ is the ratio of their fluxes. The terms omitted in (6) are proportional to $\cos\phi$ or $\cos 2\phi$ where $\phi$ is the azimuthal angle of $\bm{k}$} relative to the lepton scattering plane. These terms arise from interference between different photon polarization states. We neglect them assuming an implicit integration over $\phi_{k}$. In Ultraperipheral Collisions (UPCs), only the transverse part is relevant
$\frac{d\sigma^{\rm UPC}}{dzd^{2}{\bm{k}}d^{2}{\bm{\Delta}}}=\int dq^{0}\frac{dN}{dq^{0}}\frac{d\sigma^{T}}{dzd^{2}{\bm{k}}d^{2}{\bm{\Delta}}},$ (7)
where $dN/dq^{0}$ is the well-known equivalent photon flux Jackson:1998nia.
In Braun:2005rg, $d\sigma^{L/T}$ has been calculated in collinear factorization using the GPDs. The result reads
$\frac{d\sigma^{L}}{dzd^{2}\boldsymbol{k}d^{2}{\bm{\Delta}}}$ $=$ $\frac{\alpha_{em}e_{q}^{2}z^{2}\bar{z}^{2}Q^{2}\alpha_{s}^{2}}{2\pi^{2}N_{c}(1-\xi^{2})(k_{\perp}^{2}+\mu^{2})^{4}}|I_{L}^{g}+2C_{F}I_{L}^{q}|^{2}$ (8)
$\frac{d\sigma^{T}}{dzd^{2}\boldsymbol{k}d^{2}{\bm{\Delta}}}$ $=$ $\frac{\alpha_{em}e_{q}^{2}\alpha_{s}^{2}}{16\pi^{2}N_{c}(1-\xi^{2})(k_{\perp}^{2}+\mu^{2})^{4}}$ (9)
$\quad\times\biggl(k_{\perp}^{2}\Bigl(\bigl|2C_{F}(I_{T}^{q_{1}}+I_{T}^{q_{2}})-(1-2z)I_{T}^{g}\bigr|^{2}+\bigl|2C_{F}(I_{T}^{q_{1}}-I_{T}^{q_{2}})+I_{T}^{g}\bigr|^{2}\Bigr)+2m^{2}|I_{L}^{g}|^{2}\biggr),$
where $\mu^{2}=m^{2}+z\bar{z}Q^{2}$ and
$I_{L}^{g}$ $=$ $\int_{-1}^{1}dxF_{g}(x,\xi)\left(\frac{4\xi(1-\beta)(x^{2}+\xi^{2})}{(x^{2}-\xi^{2}+i\epsilon)^{2}}-\frac{2\xi(1-2\beta)}{x^{2}-\xi^{2}+i\epsilon}\right)$
$I_{L}^{q}$ $=$ $\int_{-1}^{1}dxF_{q}(x,\xi)\left(\frac{2\xi\bar{z}}{x+\xi-i\epsilon}+\frac{2\xi z}{x-\xi+i\epsilon}\right),$
$I_{T}^{g}$ $=$ $\int_{-1}^{1}dxF_{g}(x,\xi)\left(\frac{2\xi(1-2\beta)(x^{2}+\xi^{2})}{(x^{2}-\xi^{2}+i\epsilon)^{2}}+\frac{4\xi\beta}{x^{2}-\xi^{2}+i\epsilon}\right),$ (10)
$I_{T}^{q_{1}}$ $=$ $\int_{-1}^{1}dxF_{q}(x,\xi)\left(\frac{2\xi z\bar{z}}{x-\xi+i\epsilon}-\frac{2\xi\beta\bar{z}^{2}}{(1-\beta)(x+\xi-i\epsilon)}+\frac{2\xi\bar{z}^{2}}{(1-\beta)(x-\xi(1-2\beta)-i\epsilon)}\right),$
$I_{T}^{q_{2}}$ $=$ $\int_{-1}^{1}dxF_{q}(x,\xi)\left(\frac{2\xi\beta z^{2}}{(1-\beta)(x-\xi+i\epsilon)}-\frac{2\xi z\bar{z}}{x+\xi-i\epsilon}-\frac{2\xi z^{2}}{(1-\beta)(x+\xi(1-2\beta)+i\epsilon)}\right).$
$I_{L}^{q},I_{T}^{q_{1,2}}$ and $I_{L}^{g},I_{T}^{g}$ come from the $t$-channel quark-antiquark and two-gluon exchange diagrams, respectively. In (10), we introduced the variable
$\beta=\frac{\mu^{2}}{k_{\perp}^{2}+\mu^{2}},$ (11)
often used in diffractive DIS and defined
$F_{q,g}\equiv H_{q,g}-\frac{\xi^{2}}{1-\xi^{2}}E_{q,g}.$ (12)
The quark and gluon GPDs $H_{q}$ and $H_{g}$ are normalized such that they reduce to the respective PDFs $H_{q}(x,0)=q(x)$ and $H_{g}(x,0)=xg(x)$ in the forward limit. In this paper we are only concerned with the region $\xi\ll 1$, and accordingly neglect ${\cal O}(\xi^{2})$ terms. The GPD $E$’s drop out in this approximation.
Figure 1: $q\bar{q}$ production in the photon-Pomeron CM frame at $\phi=0$. Left: $z>\frac{1}{2}$, Right: $z<\frac{1}{2}$.
II.1 Longitudinal photon
The cross sections (8), (9) have been obtained by summing over the two spin states of the produced quark and antiquark. In order to compute the spin density matrix, one has to undo this step. Let $\alpha,\alpha^{\prime}$ be the spin indices of the quark and the antiquark in the amplitude, and $\beta,\beta^{\prime}$ be those in the complex-conjugate amplitude. Before setting $\alpha=\beta$, $\alpha^{\prime}=\beta^{\prime}$ and summing over $\alpha,\alpha^{\prime}$, the squared amplitude in the longitudinal case is proportional to Braun:2005rg
[…]
This is exactly the same as in the Pomeron exchange contribution calculated in Fucilla:2025kit. We then immediately see that the spin density matrix $C_{ab}^{L}$ is the same. It came as a surprise in Fucilla:2025kit that $C_{ab}^{L}$ resulting from the Pomeron exchange is identical to that for the lowest order process $\gamma^{*}_{L}+g\to q+\bar{q}$ calculated in Qi:2025onf. That is, $C_{ab}^{L}$ does not depend on whether color singlet or octet states are exchanged in the $t$-channel, or whether one gluon or multiple gluons are exchanged. We now see that $C^{L}_{ab}$ remains the same even after including the quark-antiquark exchange channel.
For completeness, and to set up our notation for the transversely polarized case to be discussed later, here we reproduce the result. First we go to the photon-Pomeron CM frame where the quark and the antiquark have back-to-back momenta (switching to the Cartesian coordinates $k^{\mu}=(k^{0},\vec{k})=(k^{0},{\bm{k}},k^{3})$)
$k^{\mu}=\left(\frac{M}{2},{\bm{k}},\frac{M}{2}(z-\bar{z})\right)\equiv(k^{0},\vec{k}),\qquad k^{\prime\mu}=\left(\frac{M}{2},-{\bm{k}},-\frac{M}{2}(z-\bar{z})\right)=(k^{0},-\vec{k}),$ (14)
and velocity
$|\vec{k}|=\frac{M}{2}v,\qquad v=\sqrt{1-\frac{4m^{2}}{M^{2}}}.$ (15)
This can be achieved via a Lorentz boost along the $x^{3}$ axis $\tilde{k}^{\pm}=e^{\pm\eta}k^{\pm}$, $\tilde{k}^{\prime\pm}=e^{\pm\eta}k^{\prime\pm}$ with
$e^{\eta}=q^{+}\sqrt{\frac{2z\bar{z}}{k_{\perp}^{2}+m^{2}}}.$ (16)
In this frame, the quark moves in the direction $\hat{k}=\frac{\vec{k}}{|\vec{k}|}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$ with polar angles
$\cos\theta=\frac{(z-\bar{z})M}{\sqrt{M^{2}-4m^{2}}},\qquad\phi={\rm arg}(k^{1}+ik^{2}).$ (17)
Note that $\cos\theta\to 1$ when $z\to 1$ and $\cos\theta\to-1$ when $z\to 0$. Under this boost, the spinor product transforms as
$\bar{u}_{\alpha}(\tilde{k})\gamma^{+}v_{\alpha^{\prime}}(\tilde{k}^{\prime})=e^{\eta}\bar{u}_{\alpha}(k)\gamma^{+}v_{\alpha^{\prime}}(k^{\prime}).$ (18)
[…]
$\hat{n}=\frac{\hat{x}^{3}\times\hat{k}}{\sin\theta}=(-\sin\phi,\cos\phi,0),\qquad\hat{r}=\hat{k}\times\hat{n}=(-\cos\theta\cos\phi,-\cos\theta\sin\phi,\sin\theta).$ (19)
[…]
$\frac{d\sigma^{L}}{dzd^{2}{\bm{k}}d^{2}{\bm{\Delta}}}=\sum_{\alpha=\beta,\alpha^{\prime}=\beta^{\prime}}A^{L}\xi^{\dagger}_{\alpha}\eta^{\dagger}_{\alpha^{\prime}}\rho^{L}_{\alpha\beta,\alpha^{\prime}\beta^{\prime}}\xi_{\beta}\eta_{\beta^{\prime}}$ (20)
where $\xi,\eta$ are the two-component quark and antiquark spinors and
$A^{L}=\frac{\alpha_{em}e_{q}^{2}\alpha_{s}^{2}z^{2}\bar{z}^{2}Q^{2}}{2\pi^{2}N_{c}(1-\xi^{2})(k_{\perp}^{2}+\mu^{2})^{4}}|I_{L}^{g}+2C_{F}I_{L}^{q}|^{2},$ (21)
is the unpolarized cross section (8). The density matrix is most generally parametrized by
$\rho^{L}=\frac{1}{4}\left({\mathds{1}}\otimes{\mathds{1}}+B_{a}^{L}\sigma^{a}\otimes{\mathds{1}}+\bar{B}^{L}_{b}{\mathds{1}}\otimes\sigma^{b}+C_{ab}^{L}\sigma^{a}\otimes\sigma^{b}\right),\qquad(a,b=n,r,k)$ (22)
where $\mathds{1}$} is the unit $2\times 2$ matrix and $\sigma^{a}$ are the Pauli matrices. The vectors $B^{L}_{a}$, $\bar{B}^{L}_{b}$ represent the polarization of the quark and the antiquark,
$B^{L}_{a}={\rm Tr}[(\sigma^{a}\otimes{\mathds{1}})\rho^{L}],\qquad\bar{B}_{b}^{L}={\rm Tr}[({\mathds{1}}\otimes\sigma^{b})\rho^{L}],$ (23)
while the matrix $C^{L}_{ab}$ describes the correlation between the quark and antiquark spins. We find $B^{L}_{a}=\bar{B}_{b}^{L}=0$ and
$C_{nn}^{L}=1,$ $C_{rr}^{L}=-C_{kk}^{L}=-\frac{1-(2-v^{2})\cos^{2}\theta}{1-v^{2}\cos^{2}\theta}=-\frac{k_{\perp}^{2}-(1-2z)^{2}m^{2}}{k_{\perp}^{2}+(1-2z)^{2}m^{2}},$ (24)
$C_{rk}^{L}=C_{kr}^{L}=-\frac{\sqrt{1-v^{2}}\sin 2\theta}{1-v^{2}\cos^{2}\theta}=\frac{2(1-2z)k_{\perp}m}{k_{\perp}^{2}+(1-2z)^{2}m^{2}}.$
The other components are zero. Note that $(C_{rr}^{L})^{2}+(C_{rk}^{L})^{2}=1$.
The matrix $C_{ab}^{L}$ describes a two-qubit state that is maximally entangled and moreover maximally violates the Bell-CHSH inequality Qi:2025onf. In particular, for the symmetric pair $z=\frac{1}{2}$ ($\cos\theta=0$), or in the relativistic limit $k_{\perp}\to\infty$, or in the massless case $m=0$, $C^{L}={\rm diag}(1,-1,1)$ represents one of the Bell states
$\rho^{L}=|\Phi^{+}\rangle\langle\Phi^{+}|,\qquad|\Phi^{+}\rangle=\frac{1}{\sqrt{2}}\bigl(|++\rangle+|--\rangle\bigr).$ (25)
where $\pm$ refers to spin projection $\pm\frac{1}{2}$ along the $\hat{k}-$axis. (This is opposite in sign to helicity for the antiquark moving in the $\hat{k}$ direction.) On the other hand, in the forward limit $k_{\perp}\to 0$ ($\cos\theta=\pm 1$), $C^{L}={\rm diag}(1,1,-1)$ represents another Bell state
$\rho^{L}=|\Psi^{+}\rangle\langle\Psi^{+}|,\qquad|\Psi^{+}\rangle=\frac{1}{\sqrt{2}}\bigl(|+-\rangle+|-+\rangle\bigr).$ (26)
Moreover, along the line $z=\frac{1}{2}\pm\frac{k_{\perp}}{2m}$,
$C^{L}=\begin{pmatrix}1&0&0\\ 0&0&\mp 1\\ 0&\mp 1&0\end{pmatrix}$ (27)
This represents a state equivalent to the Bell state up to a local phase rotation
$\rho^{L}=|\Phi_{n}^{\pm}\rangle\langle\Phi_{n}^{\pm}|,\qquad|\Phi_{n}^{\pm}\rangle\equiv\frac{1}{\sqrt{2}}\left(|++\rangle_{n}\pm i|--\rangle_{n}\right),$ (28)
where the subscript $n$ means that the spin eigenstates are defined with respect to the $\hat{n}$ axis.
II.2 Transverse photon
We now turn to the transversely polarized case that has a much richer structure from the viewpoint of entanglement. The amplitude is proportional to Braun:2005rg
[…]
$=e^{\eta}\epsilon^{i}\bar{u}_{\alpha}(k)\Biggl[-m\gamma^{i}\gamma^{+}I_{L}^{g}+k_{\perp}^{i}\gamma^{+}X+i\epsilon^{ij}k_{\perp j}\gamma^{+}\gamma_{5}Y\Biggr]v_{\alpha^{\prime}}(k^{\prime}),$ (29)
where we defined
$X=2C_{F}(I_{T}^{q_{1}}+I_{T}^{q_{2}})+(\bar{z}-z)I_{T}^{g},\qquad Y=2C_{F}(I_{T}^{q_{1}}-I_{T}^{q_{2}})+I_{T}^{g}.$ (30)
We used the fact that $\gamma^{i=1,2}$ and $\gamma_{5}$ are invariant under the boost. Since we focus on the $\phi$-independent part of the cross section, we may set $\phi=0$ without losing generality. Squaring (29) and summing over the polarizations of the incoming photon $\epsilon^{i}\epsilon^{j*}\to\delta^{ij}$, we find that the cross section is proportional to
$k_{\perp}^{2}|X|^{2}\bar{u}_{\alpha}(k)\gamma^{+}v_{\alpha^{\prime}}(k^{\prime})\bar{v}_{\beta^{\prime}}(k^{\prime})\gamma^{+}u_{\beta}(k)+k_{\perp}^{2}|Y|^{2}\bar{u}_{\alpha}(k)\gamma^{+}\gamma_{5}v_{\alpha^{\prime}}(k^{\prime})\bar{v}_{\beta^{\prime}}(k^{\prime})\gamma^{+}\gamma_{5}u_{\beta}(k)$
$-m^{2}|I_{L}^{g}|^{2}\bar{u}_{\alpha}(k)\gamma^{+}\gamma^{i}v_{\alpha^{\prime}}(k^{\prime})\bar{v}_{\beta^{\prime}}(k^{\prime})\gamma^{+}\gamma^{i}u_{\beta}(k)$
$-mk_{\perp}\bigg[XI_{L}^{g*}\bar{u}_{\alpha}(k)\gamma^{+}v_{\alpha^{\prime}}(k^{\prime})\bar{v}_{\beta^{\prime}}(k^{\prime})\gamma^{+}\gamma^{1}u_{\beta}(k)+X^{*}I_{L}^{g}\bar{u}_{\alpha}(k)\gamma^{1}\gamma^{+}v_{\alpha^{\prime}}(k^{\prime})\bar{v}_{\beta^{\prime}}(k^{\prime})\gamma^{+}u_{\beta}(k)\bigg]$
$-imk_{\perp}\bigg[YI_{L}^{g*}\bar{u}_{\alpha}(k)\gamma^{+}\gamma^{5}v_{\alpha^{\prime}}(k^{\prime})\bar{v}_{\beta^{\prime}}(k^{\prime})\gamma^{+}\gamma^{2}u_{\beta}(k)+Y^{*}I_{L}^{g}\bar{u}_{\alpha}(k)\gamma^{+}\gamma^{2}v_{\alpha^{\prime}}(k^{\prime})\bar{v}_{\beta^{\prime}}(k^{\prime})\gamma^{+}\gamma^{5}u_{\beta}(k)\bigg].$ (31)
[…]
$\gamma^{+}=\frac{1}{\sqrt{2}}\left(\gamma^{0}+\sin\theta\gamma^{r}+\cos\theta\gamma^{k}\right),\qquad\gamma^{1}=-\cos\theta\gamma^{r}+\sin\theta\gamma^{k},\qquad\gamma^{2}=\gamma^{n}.$ (32)
We then express the cross section in the form (20), (22) with $L\to T$. The spin-averaged part is as in (9),
$A^{T}$ $=$ $\frac{\alpha_{em}e_{q}^{2}\alpha_{s}^{2}}{16\pi^{2}N_{c}(1-\xi^{2})(k_{\perp}^{2}+\mu^{2})^{4}}\left(k_{\perp}^{2}(|X|^{2}+|Y|^{2})+2m^{2}|I_{L}^{g}|^{2}\right).$ (33)
For the polarization vectors, we find $B^{T}_{r,k}=\bar{B}^{T}_{r,k}=0$, but the $n$-component turns out to be nonvanishing
$B_{n}^{T}=\frac{-2mk_{\perp}{\rm Im}[(X+Y)I_{L}^{g*}]}{k_{\perp}^{2}(|X|^{2}+|Y|^{2})+2m^{2}|I_{L}^{g}|^{2}},\qquad\bar{B}_{n}^{T}=\frac{-2mk_{\perp}{\rm Im}[(X-Y)I_{L}^{g*}]}{k_{\perp}^{2}(|X|^{2}+|Y|^{2})+2m^{2}|I_{L}^{g}|^{2}}.$ (34)
This means that the quark and the antiquark are transversely (i.e., normal to the production plane, along the $\hat{n}-$direction in Fig. 1) polarized. It can be readily shown that
$|B_{n}^{T}|,|\bar{B}_{n}^{T}|\leq\frac{\left|{\rm Im}[(X\pm Y)I_{L}^{g*}]\right|}{\sqrt{2|I_{L}^{g*}|^{2}(|X|^{2}+|Y|^{2})}}\leq\frac{\left|X\pm Y\right|}{\sqrt{2(|X|^{2}+|Y|^{2})}}\leq 1,$ (35)
where in the first inequality we maximized in $k_{\perp}$. The last inequality is saturated when $X=Y$ and X=YX=-Y, which occurs at $z=0$ and $\bar{z}=0$, respectively. Since the colliding particles $e+p$ or $\gamma+p$ are unpolarized, (34) can be interpreted as single spin asymmetry (SSA), analogous to the transverse polarization of hyperons in unpolarized hadron collisions Bunce:1976yb; Lundberg:1989hw; HERMES:2007fpi; COMPASS:2021bws. [^3] The transverse polarization of final state particles can also arise if initial state particles are transversely polarized. See Cheng:2025zaw for a recent study at the EIC. As is well known, SSA requires a phase, and therefore it cannot be generated in tree-level calculations in parity-conserving theories like QCD. Theoretical descriptions usually invoke higher-twist distribution and/or fragmentation functions, or their transverse momentum dependent generalizations Liang:1997rt; Kanazawa:2000cx; Anselmino:2001js; Zhou:2008fb; Koike:2017fxr; Kang:2021kpt. However, a phase can also be generated in perturbation theory beyond the tree level Dharmaratna:1996xd; Benic:2024fvk, and this is what happens here. Since GPD amplitudes involve a loop, they are naturally complex. (Notice the $i\epsilon$ terms in (10).) (34) clearly shows that the nonvanishing polarization results from the interference between their real and imaginary parts. In contrast, in the Regge limit, $X,Y$ and $I_{L}^{g}$ become purely imaginary, see the next subsection. This is why no polarization was found in Fucilla:2025kit.
As for the $C^{T}$ matrix, we obtain
$C^{T}_{nn}=\frac{(|X|^{2}-|Y|^{2})k_{\perp}^{2}}{k_{\perp}^{2}(|X|^{2}+|Y|^{2})+2m^{2}|I_{L}^{g}|^{2}},$ (36)
$C^{T}_{rr}$ $=$ $\frac{k_{\perp}^{2}}{k_{\perp}^{2}(|X|^{2}+|Y|^{2})+2m^{2}|I_{L}^{g}|^{2}}\Biggl[-\frac{k_{\perp}^{2}-(1-2z)^{2}m^{2}}{k_{\perp}^{2}+(1-2z)^{2}m^{2}}|X|^{2}+\frac{k_{\perp}^{2}-m^{2}}{k_{\perp}^{2}+m^{2}}|Y|^{2}$
$\qquad+\frac{8z\bar{z}m^{4}|I_{L}^{g}|^{2}}{(k_{\perp}^{2}+(1-2z)^{2}m^{2})(k_{\perp}^{2}+m^{2})}+4m^{2}\left(-\frac{(1-2z){\rm Re}[XI_{L}^{g*}]}{k_{\perp}^{2}+(1-2z)^{2}m^{2}}+\frac{{\rm Re}[YI_{L}^{g*}]}{k_{\perp}^{2}+m^{2}}\right)\Biggr],$
$C^{T}_{kk}$ $=$ $\frac{1}{k_{\perp}^{2}(|X|^{2}+|Y|^{2})+2m^{2}|I_{L}^{g}|^{2}}\Biggl[k_{\perp}^{2}\frac{k_{\perp}^{2}-(1-2z)^{2}m^{2}}{k_{\perp}^{2}+(1-2z)^{2}m^{2}}|X|^{2}+k_{\perp}^{2}\frac{k_{\perp}^{2}-m^{2}}{k_{\perp}^{2}+m^{2}}|Y|^{2}$
$\qquad-\frac{2m^{2}(k_{\perp}^{4}-(1-2z)^{2}m^{4})|I_{L}^{g}|^{2}}{(k_{\perp}^{2}+(1-2z)^{2}m^{2})(k_{\perp}^{2}+m^{2})}+4m^{2}k_{\perp}^{2}\left(\frac{(1-2z){\rm Re}[XI_{L}^{g*}]}{k_{\perp}^{2}+(1-2z)^{2}m^{2}}+\frac{{\rm Re}[YI_{L}^{g*}]}{k_{\perp}^{2}+m^{2}}\right)\Biggr],$
$C^{T}_{rk}$ $=$ $\frac{2k_{\perp}m}{k_{\perp}^{2}(|X|^{2}+|Y|^{2})+2m^{2}|I_{L}^{g}|^{2}}\Biggl[\frac{(1-2z)k_{\perp}^{2}|X|^{2}}{k_{\perp}^{2}+(1-2z)^{2}m^{2}}+\frac{k_{\perp}^{2}|Y|^{2}}{k_{\perp}^{2}+m^{2}}$
$\quad-\frac{2(1-z)m^{2}(k_{\perp}^{2}+(1-2z)m^{2})|I_{L}^{g}|^{2}}{(k_{\perp}^{2}+m^{2})(k_{\perp}^{2}+(1-2z)^{2}m^{2})}-\frac{k_{\perp}^{2}-(1-2z)^{2}m^{2}}{k_{\perp}^{2}+(1-2z)^{2}m^{2}}{\rm Re}[XI_{L}^{g*}]-\frac{k_{\perp}^{2}-m^{2}}{k_{\perp}^{2}+m^{2}}{\rm Re}[YI_{L}^{g*}]\Biggr],$
$C^{T}_{kr}$ $=$ $\frac{2k_{\perp}m}{k_{\perp}^{2}(|X|^{2}+|Y|^{2})+2m^{2}|I_{L}^{g}|^{2}}\Biggl[\frac{(1-2z)k_{\perp}^{2}|X|^{2}}{k_{\perp}^{2}+(1-2z)^{2}m^{2}}-\frac{k_{\perp}^{2}|Y|^{2}}{k_{\perp}^{2}+m^{2}}$
$\qquad+\frac{2zm^{2}(k_{\perp}^{2}-(1-2z)m^{2})|I_{L}^{g}|^{2}}{(k_{\perp}^{2}+m^{2})(k_{\perp}^{2}+(1-2z)^{2}m^{2})}-\frac{k_{\perp}^{2}-(1-2z)^{2}m^{2}}{k_{\perp}^{2}+(1-2z)^{2}m^{2}}{\rm Re}[XI_{L}^{g*}]+\frac{k_{\perp}^{2}-m^{2}}{k_{\perp}^{2}+m^{2}}{\rm Re}[YI_{L}^{g*}]\Biggr].$
Unlike in the longitudinal case, the $C^{T}$ matrix depends on the nonperturbative structure of the proton through their dependence on the quark and gluon GPDs.
One might have expected that $B^{T}_{n}(z)=\bar{B}^{T}_{n}(z)$ and $C^{T}_{rk}(z)=C^{T}_{kr}(z)$ from CP symmetry, as in the case of $q\bar{q}\to t\bar{t}$ and $gg\to t\bar{t}$ productions in QCD Bernreuther:2015yna. However, neither of these relations hold in the present problem since the initial state $\gamma+{\mathbb{P}}$ is asymmetric (see also Qi:2025onf). Instead, the following relations hold exactly
$B^{T}_{n}(z)=-\bar{B}^{T}_{n}(\bar{z}),\qquad C^{T}_{rk}(z)=-C^{T}_{kr}(\bar{z}).$ (41)
For the valence quarks which we treat as massless $m_{u,d}=0$, this is trivial. For massive quarks $q=s,c,b$, (41) is a consequence of the property $F_{q}(x,\xi)=-F_{q}(-x,\xi)$ of the nonvalence quark GPDs which results in the relations
$X(z)=-X(\bar{z}),\qquad Y(z)=Y(\bar{z}).$ (42)
We may regard (41) as constraints from CP symmetry, since $z\to\bar{z}$ (or $\cos\theta\to-\cos\theta$ in the CM frame) interchanges the quark and the antiquark.
II.3 Regge limit
Let us quickly check consistency with the result of Fucilla:2025kit obtained in the Regge limit $\xi\to 0$. In this limit, the quark exchange contributions $I_{L}^{q},I_{T}^{q_{1}},I_{T}^{q_{2}}$ are subdominant and neglected altogether. The gluon exchange contributions $I_{L}^{g},I_{T}^{g}$ are dominantly imaginary, coming from the $i\epsilon$ prescription in the integrals (10). Keeping only the most singular term in the imaginary part, one finds Braun:2005rg
$I_{L}^{g}\approx 2\pi i\frac{k_{\perp}^{2}-\mu^{2}}{k_{\perp}^{2}+\mu^{2}}H_{g}(\xi,\xi),\qquad I_{T}^{g}\approx 2i\pi\frac{-2\mu^{2}}{k_{\perp}^{2}+\mu^{2}}H_{g}(\xi,\xi),$ (43)
where it is assumed that the gluon GPD has a Regge behavior $H_{g}(\xi,\xi)\sim 1/\xi^{\alpha}$ with $\alpha\ll 1$. On the other hand, the result of Fucilla:2025kit is expressed in terms of the following two integrals
$T_{1}=\int\frac{d^{2}{\bm{p}}T(p)}{({\bm{p}}-{\bm{k}})^{2}+\mu^{2}},\qquad T_{2}=-\frac{1}{k_{\perp}^{2}}\int d^{2}p\frac{{\bm{p}}\cdot{\bm{k}}T(p)}{({\bm{p}}-{\bm{k}})^{2}+\mu^{2}},$ (44)
where $T$ is the T-matrix of the $q\bar{q}$ pair (‘color dipole’) scattering off the target proton. To make a connection with the collinear GPD calculation, we expand the integrand in powers of the intrinsic transverse momentum $\bm{p}$} and use the relation between $T(p)$ and $H_{g}(\xi,\xi)$ Hatta:2017cte. We find
$T_{1}$ $\approx$ $\frac{k_{\perp}^{2}-\mu^{2}}{(k_{\perp}^{2}+\mu^{2})^{3}}\int d^{2}{\bm{p}}\,{\bm{p}}^{2}T(p)\approx\frac{k_{\perp}^{2}-\mu^{2}}{2(k_{\perp}^{2}+\mu^{2})^{3}}\alpha_{s}H_{g}(\xi,\xi),$
$T_{2}$ $\approx$ $-\frac{1}{(k_{\perp}^{2}+\mu^{2})^{2}}\int d^{2}{\bm{p}}\,{\bm{p}}^{2}T(p)\approx-\frac{1}{2(k_{\perp}^{2}+\mu^{2})^{2}}\alpha_{s}H_{g}(\xi,\xi).$ (45)
This gives the correspondence
$T_{1}\approx\frac{\alpha_{s}I_{L}^{g}}{4\pi i(k_{\perp}^{2}+\mu^{2})^{2}},\qquad T_{1}+T_{2}\approx\frac{\alpha_{s}I_{T}^{g}}{4\pi i(k_{\perp}^{2}+\mu^{2})^{2}}.$ (46)
Inserting these relations into $A^{T/L}$, and noting that $X\approx(\bar{z}-z)I_{T}^{g}$, $Y\approx I_{T}^{g}$, one recovers the dijet cross section originally calculated in the $k_{\perp}$-factorization framework Nikolaev:1994cd; Bartels:1996ne, as already noted in Braun:2005rg. Furthermore, inserting into (36)-(II.2), we find full agreement with Fucilla:2025kit. Note that $T_{1}$ linearly vanishes as $k_{\perp}\to\mu$ Fucilla:2025kit. Consistently, $I_{L}^{g}$ has the same behavior in the present approximation.
III Entanglement, Bell nonlocality and magic
In this section, we discuss three different measures of spin-spin correlation between the produced quark and antiquark: entanglement, Bell nonlocality and magic. The spin density matrix formalism allows one to apply these key concepts in quantum information science to relativistic field theories. Numerical results will be presented in the next section.
III.1 Entanglement
According to the Peres-Hordecki criterion Peres:1996dw; Horodecki:1997vt, for generic two-qubit systems, the necessary and sufficient condition for separability (i.e., no entanglement) is that the partial transpose of the density matrix
$\rho^{\cal T}\equiv\frac{1}{4}\left({\mathds{1}}\otimes{\mathds{1}}^{\cal T}+B_{a}\sigma^{a}\otimes{\mathds{1}}^{\cal T}+\bar{B}_{b}{\mathds{1}}\otimes(\sigma^{b})^{\cal T}+C_{ab}\sigma^{a}\otimes(\sigma^{b})^{\cal T}\right),$ (47)
is nonnegative definite. (The symbol $\cal T$} denotes ‘transpose.’) In the present problem with $\rho=\rho^{L/T}$, it is convenient to use the spin eigenstates along the direction $\hat{n}$. In this basis, the density matrix is block diagonal Afik:2020onf
$\hskip-8.53581pt\rho^{\cal T}=$ (48)
$\hskip-8.53581pt\frac{1}{4}\begin{pmatrix}1+B_{n}+\bar{B}_{n}+C_{nn}&0&0&C_{rr}+C_{kk}+i(C_{rk}-C_{kr})\\ 0&1+B_{n}-\bar{B}_{n}-C_{nn}&C_{rr}-C_{kk}-i(C_{rk}+C_{kr})&0\\ 0&C_{rr}-C_{kk}+i(C_{rk}+C_{kr})&1-B_{n}+\bar{B}_{n}-C_{nn}&0\\ C_{rr}+C_{kk}-i(C_{rk}-C_{kr})&0&0&1-B_{n}-\bar{B}_{n}+C_{nn}\end{pmatrix}$.
In order for this matrix to be nonnegative, the following conditions need to be satisfied
$1\pm C_{nn}\geq 0,\qquad(1\pm C_{nn})^{2}-(B_{n}\pm\bar{B}_{n})^{2}-(C_{rr}\pm C_{kk})^{2}-(C_{rk}\mp C_{kr})^{2}\geq 0.$ (49)
Thus the necessary and sufficient condition for entanglement is
${\rm max}\{\Delta_{1},\Delta_{2},|C_{nn}|-1\}>0,$ (50)
where
$\Delta_{1}\equiv\sqrt{(C_{rr}-C_{kk})^{2}+(C_{rk}+C_{kr})^{2}+(B_{n}-\bar{B}_{n})^{2}}-|1-C_{nn}|,$
$\Delta_{2}\equiv\sqrt{(C_{rr}+C_{kk})^{2}+(C_{rk}-C_{kr})^{2}+(B_{n}+\bar{B}_{n})^{2}}-|1+C_{nn}|.$ (51)
In Fucilla:2025kit, it has been noticed that in the Regge limit,
$\Delta^{L/T}_{2}=-\Delta^{L/T}_{1}=-2C^{L/T}_{nn}.$ (52)
In the longitudinal case, the $q\bar{q}$ pair is always maximally entangled because $C_{nn}^{L}=1$ and $\Delta_{1}^{L}=2>0$. In the transverse case where $\Delta_{2}^{T}=-2C_{nn}^{T}\geq 0$, the pair is almost always entangled except when $C_{nn}^{T}=0$, which occurs in certain kinematical limits such as $k_{\perp}=0$ or $k_{\perp}=\infty$. (52) is a direct consequence of the following relations obeyed by the components of $C^{L/T}$
$(C_{rr})^{2}+(C_{rk})^{2}+(C_{kr})^{2}+(C_{kk})^{2}-(C_{nn})^{2}-1=0,\qquad C_{nn}+C_{rr}C_{kk}-C_{rk}C_{kr}=0.$ (53)
Away from the Regge limit, (53) continue to hold in the longitudinal case. However, in the transverse case we find from (36)-(II.2)
$\begin{split}&(C_{rr}^{T})^{2}+(C_{rk}^{T})^{2}+(C_{kr}^{T})^{2}+(C_{kk}^{T})^{2}-(C_{nn}^{T})^{2}-1\\ &\qquad=8k_{\perp}^{2}m^{2}\frac{({\rm Re}[XI_{L}^{g*}])^{2}+({\rm Re}[YI_{L}^{g*}])^{2}-(|X|^{2}+|Y|^{2})|I_{L}^{g}|^{2}}{((|X|^{2}+|Y|^{2})k_{\perp}^{2}+2|I_{L}^{g}|^{2}m^{2})^{2}},\\ &C_{nn}^{T}+C_{rr}^{T}C_{kk}^{T}-C_{rk}^{T}C_{kr}^{T}=4k_{\perp}^{2}m^{2}\frac{-({\rm Re}[XI_{L}^{g*}])^{2}+({\rm Re}[YI_{L}^{g*}])^{2}+(|X|^{2}-|Y|^{2})|I_{L}^{g}|^{2}}{((|X|^{2}+|Y|^{2})k_{\perp}^{2}+2|I_{L}^{g}|^{2}m^{2})^{2}}.\end{split}$ (54)
In the Regge limit, the right hand sides vanish because $X,Y,I_{L}^{g}$ are purely imaginary. However, in general $X,Y,I_{L}^{g}$ have both real and imaginary parts. [^4] The right hand sides also vanish if $X,Y,I_{L}^{g}$ are all relatively real. Again, this not the case in general. Therefore, in the present calculation (52) does not hold, and the condition (50) becomes nontrivial to satisfy. We expect to find regions in phase space where the $q\bar{q}$ pair is separable, i.e., not entangled. Interestingly, the right hand sides of (54) vanish for massless quarks $m=0$. This means that the $u\bar{u}$ and $d\bar{d}$ pairs, and also the $s\bar{s}$ pair in practice, are always entangled.
Let us elaborate more on the massless case. The $C^{T}$ matrix becomes diagonal
$C^{T}_{ab}=\begin{pmatrix}\frac{|X|^{2}-|Y|^{2}}{|X|^{2}+|Y|^{2}}&0&0\\ 0&-\frac{|X|^{2}-|Y|^{2}}{|X|^{2}+|Y|^{2}}&0\\ 0&0&1\end{pmatrix}$. (55)
The corresponding density matrix can be written in the form
$\rho^{T}=\frac{|X|^{2}}{|X|^{2}+|Y|^{2}}|\Phi^{+}\rangle\langle\Phi^{+}|+\frac{|Y|^{2}}{|X|^{2}+|Y|^{2}}|\Phi^{-}\rangle\langle\Phi^{-}|.$ (56)
This represents a mixed state between two of the Bell states.
$|\Phi^{\pm}\rangle=\frac{1}{\sqrt{2}}\bigl(|++\rangle\pm|--\rangle\bigr).$ (57)
(56) is an example of the so-called ‘Bell diagonal’ states Lang:2010xtw which are mixed states comprised only of the Bell states
$\rho=p_{1}|\Psi^{+}\rangle\langle\Psi^{+}|+p_{2}|\Psi^{-}\rangle\langle\Psi^{-}|+p_{3}|\Phi^{+}\rangle\langle\Phi^{+}|+p_{4}|\Phi^{-}\rangle\langle\Phi^{-}|,$ (58)
with $\sum_{i}p_{i}=1$. It is known that a Bell diagonal state is separable if $p_{i}\leq\frac{1}{2}$ for all $i$. In the present problem, the pair is separable only if $|X|=|Y|$ where $C_{nn}^{T}=0$. If either $|X|$ or $|Y|$ (accidentally) vanishes, then one of the Bell states is realized. Note that in the Regge limit, $X\approx(1-2z)I_{T}^{g}$ and $Y\approx I_{T}^{g}$. Thus the state is separable when $z=0,1$, and becomes the Bell state $|\Phi^{-}\rangle$ at $z=\frac{1}{2}$.
III.2 Bell nonlocality
In terms of the correlation matrix $C$, the Bell-CHSH inequality reads Bell:1964kc; Clauser:1969ny
${\rm Max}_{\{\vec{n}_{i}\}}\Bigl|n_{1}^{a}C_{ab}(n_{2}^{b}+n_{4}^{b})+n_{3}^{a}C_{ab}(n_{2}^{b}-n_{4}^{b})\Bigr|\leq 2,$ (59)
where $\vec{n}_{i}$ ($i=1,2,3,4$) are unit vectors $|\vec{n}_{i}|=1$. Depending on the matrix $C_{ab}$, quantum two-qubit states can violate this inequality for deliberate choices of $\{\vec{n}_{i}\}$. When this happens, we say the $q\bar{q}$ pair exhibits ‘Bell nonlocality.’ The necessary and sufficient condition for Bell nonlocality is that the largest two of the three eigenvalues $\mu_{3}\leq\mu_{2}\leq\mu_{1}$ of the matrix $C^{\cal T}C$ satisfy Horodecki:1995nsk
$1<\mu_{1}+\mu_{2}\leq 2.$ (60)
For our problem, the three eigenvalues of
$C^{\cal T}C=\begin{pmatrix}C_{nn}^{2}&0&0\\ 0&C_{rr}^{2}+C_{kr}^{2}&C_{rr}C_{rk}+C_{kr}C_{kk}\\ 0&C_{rr}C_{rk}+C_{kr}C_{kk}&C_{kk}^{2}+C_{rk}^{2}\end{pmatrix}$, (61)
are $C_{nn}^{2}$ and
$\frac{C_{rr}^{2}+C_{kr}^{2}+C_{kk}^{2}+C_{rk}^{2}\pm\sqrt{\left(C_{rr}^{2}+C_{kr}^{2}-C_{kk}^{2}-C_{rk}^{2}\right)^{2}+4\left(C_{rr}C_{rk}+C_{kr}C_{kk}\right)^{2}}}{2}.$ (62)
In the Regge limit, the identities (53) hold, and the above three eigenvalues reduce to
$\mu_{1}=1,\qquad\mu_{2}=\mu_{3}=C_{nn}^{2}.$ (63)
Thus the inequality is always maximally violated in the longitudinal case $C_{nn}^{L}=1$, while in the transverse case it is violated if $C_{nn}^{T}\neq 0$. Since $C_{nn}^{T}\leq 0$, this happens to be the same condition as (50), (52) for nonvanishing entanglement Fucilla:2025kit. In other words, Bell nonlocality and entanglement are equivalent. This is remarkable because, in general, the states that exhibit Bell nonlocality form a subset of the states that are entangled. Away from the Regge limit, the Bell-CHSH inequality continues to be maximally violated in the longitudinal case. In the transverse case, it is always violated for the $q=u,d$ quarks because (53) also holds for massless quarks. For massive quarks, there will be kinematical regions where the inequality is not violated. On general grounds, we expect that this region is broader than the region where the pair is separable (not entangled).
[…]
Figure 2: Stabilizer Rényi entropy $M_{2}^{L}$ (65) of a massive quark-antiquark pair produced by a longitudinally polarized virtual photon. $M_{2}^{L}$ is suppressed in the blue regions where the pair forms the Bell states (25), (26), (28).
III.3 Magic
Finally, let us consider another measure of entanglement called ‘magic’. (See White:2024nuc for an introduction friendly to high energy physicists.) It is known that entanglement alone does not guarantee quantum advantage Gottesman:1998hu. Entangled pairs with nonvanishing magic are crucial for quantum computers to outperform classical ones, and there has been interest in the high energy physics community whether collider experiments can produce such pairs White:2024nuc; Liu:2025qfl; Gargalionis:2025iqs. To quantify magic, we adopt the ‘stabilizer Rényi entropy’ Leone:2021rzd; White:2024nuc which takes the following form for the present problem
$M_{2}=-\ln\left(\frac{1+B_{n}^{4}+\bar{B}_{n}^{4}+C_{nn}^{4}+C_{rr}^{4}+C_{kk}^{4}+C_{kr}^{4}+C_{rk}^{4}}{1+B_{n}^{2}+\bar{B}_{n}^{2}+C_{nn}^{2}+C_{rr}^{2}+C_{kk}^{2}+C_{kr}^{2}+C_{rk}^{2}}\right).$ (64)
Let us study the longitudinal case which allows for a simple analytical treatment, deferring the transverse case to the numerical section. From (24), we find a compact analytical formula
$M_{2}^{L}=\ln\frac{(k_{\perp}^{2}+m^{2}(1-2z)^{2})^{4}}{k_{\perp}^{8}+14k_{\perp}^{4}m^{4}(1-2z)^{4}+m^{8}(1-2z)^{8}}.$ (65)
This is plotted in Fig. 2. By construction, $M_{2}$ vanishes for the Bell states $|\Phi^{+}\rangle$ (25), $|\Psi^{+}\rangle$ (26), $|\Phi_{n}^{\pm}\rangle$ (28) realized along the lines $z=\frac{1}{2}$, $k_{\perp}=0$ and $\frac{k_{\perp}}{m}=|2z-1|$, respectively. Everywhere else, magic is nonzero even though the pair is always maximally entangled. $M_{2}^{L}$ takes the maximal value
$M_{2}^{L,{\rm max}}=\ln\frac{4}{3}\approx 0.288,$ (66)
along the following four lines in the $(z,k_{\perp})$ plane
$\frac{k_{\perp}}{m}=(\sqrt{2}\pm 1)|2z-1|,$ (67)
which are clearly visible in Fig. 2. Along these lines, the density matrix takes the form
$\rho^{L}=\frac{1}{2}\begin{pmatrix}1&0&0&\pm\frac{1\pm i}{\sqrt{2}}\\ 0&0&0&0\\ 0&0&0&0\\ \pm\frac{1\mp i}{\sqrt{2}}&0&0&1\end{pmatrix}$, (68)
using the same basis as in (48). (The sign combination is arbitrary, except that the 14 and 41 components are complex conjugate to each other.) (66) is significantly smaller than the theoretical upper bound for pure two-qubit states Liu:2025frx
$M_{2}^{\rm max}=\ln\frac{16}{7}\approx 0.827.$ (69)
However, we point out that this limit can be saturated only by states with nonzero polarization vectors, namely, $B,\bar{B}\neq 0$ in the generic parametrization (22). Since $B^{L}=\bar{B}^{L}=0$, the maximal value (69) cannot be expected. A better comparison may be the range of $M_{2}$ for the subset of states with vanishing polarization. In Appendix A, we revisit the derivation of Liu:2025frx by imposing an extra constraint $B=\bar{B}=0$ and obtain a stronger bound
$\left.M_{2}^{\rm max}\right|_{B=\bar{B}=0}=\ln\frac{9}{5}\approx 0.588,$ (70)
which is still larger than (66) by a factor of about 2. In Appendix A, we further show that (66) is the maximum value of $M_{2}$ if one restricts to an even smaller subset of states. We note that the numbers $\ln\frac{9}{5}$ and $\ln\frac{4}{3}$ have been previously encountered in Liu:2025qfl as the maximal amount of magic that can be generated from certain initial stabilizer states in QED processes.
IV Numerical results
In this section we delineate kinematical regions in the parameter space $(z,k_{\perp})$ where the $q\bar{q}$ pair exhibits entanglement and Bell nonlocality. We also compute the stabilizer Rényi entropy $M_{2}^{T}$ for the transversely polarized photon. We employ the Goloskokov-Kroll model Goloskokov:2006hr; Goloskokov:2007nt of the quark and gluon GPDs. This model has been fitted to reproduce the exclusive vector meson production data at HERA. At small skewness, the gluon GPD features the Regge behavior
$H_{g}(\xi,\xi,\mu_{h})\sim\frac{1}{\xi^{0.1+0.06N}},\qquad N=\ln\frac{\mu_{h}^{2}}{Q_{0}^{2}},$ (71)
with $Q_{0}^{2}=4\,{\rm GeV}^{2}$ and $\mu_{h}=Q$ in DIS Goloskokov:2006hr; Goloskokov:2007nt). Thus the model should produce results consistent with Fucilla:2025kit in the Regge limit. Adapting to our process, we use $\mu_{h}={\rm max}\{Q,m,k_{\perp}\}$. The quark masses are fixed as $m_{u}=m_{d}=0$, $m_{s}=0.093$ GeV, $m_{c}=1.27$ GeV and $m_{b}=4.18$ GeV. For light quark production $q=u,d,s$, we include both the quark and gluon GPD contributions. For heavy quark production $q=c,b$, we include only the gluon GPD contribution (i.e., $I^{q}_{L/T}=0$ in this case) since the $c,b$-quark GPDs are unavailable (and expected to be negligible). Using the GK model, we numerically evaluate the integrals (10) on grid points on a $50\times 50$ lattice in the $(z,k_{\perp})$ plane
$0.1\leq z\leq 0.9,\qquad 1\,{\rm GeV}\leq k_{\perp}\leq k_{\perp}^{\rm max},$ (72)
and smoothly interpolate the results. $k_{\perp}^{\rm max}$ is chosen appropriately for a given value of the $\gamma+p$ CM energy $W$. The integrals can be evaluated in a numerically stable way. For very small values of $\xi$, this is achieved by using the contour deformation technique as explained around (43) of Hatta:2025vhs.
In DIS at $Q\neq 0$, the measured cross section (6) is the linear combination of the transverse and longitudinal contributions. Unfortunately, it is difficult to experimentally separate the two contributions at the EIC since typically $\varepsilon\approx 1$ at high energy colliders (see however Klest:2025yik). This means that the density matrices must be averaged
$A^{T}\rho^{T}+\varepsilon A^{L}\rho^{L}=\frac{A^{T}+\varepsilon A^{L}}{4}\left({\mathds{1}}\otimes{\mathds{1}}+B_{n}^{\rm DIS}\sigma^{n}\otimes{\mathds{1}}+\bar{B}_{n}^{\rm DIS}{\mathds{1}}\otimes\sigma^{n}+C_{ab}^{\rm DIS}\sigma^{a}\otimes\sigma^{b}\right),$ (73)
where
$B_{n}^{\rm DIS}=\frac{A^{T}B_{n}^{T}}{A_{T}+\varepsilon A^{L}},\qquad\bar{B}_{n}^{\rm DIS}=\frac{A^{T}\bar{B}_{n}^{T}}{A_{T}+\varepsilon A^{L}},\qquad C_{ab}^{\rm DIS}=\frac{A^{T}C_{ab}^{T}+\varepsilon A^{L}C_{ab}^{L}}{A^{T}+\varepsilon A^{L}}.$ (74)
$\Delta_{1,2}$, $\mu_{1,2}$ and $M_{2}$ defined in the previous section are calculated from these coefficients in the same way. In the following numerical results, we set $\varepsilon=1$.
IV.1 Polarization
First we show the results on the polarization of the quark $B_{n}$ and the antiquark $\bar{B}_{n}$. As we noted already, $B_{n}$ and $\bar{B}_{n}$ are nonvanishing only for massive quarks and for the transversely polarized photon, and tend to be suppressed at high CM energies. Therefore, it is best to focus on low-$W$ UPC or low-$W$ DIS at low-$Q^{2}$. In Fig. 3, we plot $B_{n}$ and $\bar{B}_{n}$ for the strange (upper panels) and charm (lower panels) quarks in UPC at $W=30$ GeV. We find that $B_{n}$ is negative and $\bar{B}_{n}$ is positive for all massive quark flavors $q=s,c,b$, consistently with (41). The net polarization along the $\hat{n}$ direction $B_{n}(z)+\bar{B}_{n}(z)\propto z-\bar{z}$ vanishes at $z=\frac{1}{2}$. The peak structure in $k_{\perp}$ is as expected from the analytic formula (34), with the peak magnitude reaching 80%, 60%, 50% for strange, charm and bottom quarks, respectively. [^5] It should be mentioned that $B_{n}$ of the strange quark is largest in the low-momentum region $k_{\perp}\gtrsim 1$ GeV where, strictly speaking, the present perturbative approach is not fully justified in photoproduction $Q=0$. These peaks occur when the quark (antiquark) is ‘soft’ $z\ll 1$ ($\bar{z}\ll 1$) in the original frame (2) in which the photon is fast moving and the $q\bar{q}$ pair is produced in the forward rapidity region, cf., the comment below (35). Negative $B_{n}$ means that the quark is polarized in the direction of $\vec{q}\times\vec{k}_{q}$, see Fig. 1. Similarly, the antiquark is polarized along the direction $\vec{q}\times\vec{k}_{\bar{q}}$.
Comments are in order regarding the magnitude of the polarization. In semi-inclusive quark or antiquark production, the imaginary part of the amplitude responsible for nonvanishing polarization is suppressed by a factor of $\alpha_{s}$ since it requires a loop. Typically, one finds (sub-)percent level asymmetries in such perturbative calculations Dharmaratna:1996xd; Benic:2024fvk. However, in exclusive production with color singlet exchanges in the $t$-channel, the real and imaginary parts of the GPD amplitudes are of the same order in $\alpha_{s}$. Consequently, $B_{n},\bar{B}_{n}$ (34) are order unity instead of order $\alpha_{s}$. Still, a polarization as large as 50-80% is quite remarkable. To our knowledge, such a strong polarization of perturbative origin has not been reported in the QCD literature.
It is generally expected that the quark-level polarization is largely retained during fragmentation into heavy baryons $q\to\Lambda_{q}$, $\bar{q}\to\bar{\Lambda}_{q}$ Galanti:2015pqa. In particular, the process $s\to\Lambda\to p+\pi^{-}$ can be used as a polarimeter of the strange quark. The HERMES collaboration HERMES:2007fpi measured the polarization of $\Lambda$ and $\bar{\Lambda}$ in semi-inclusive DIS in the photoproduction region $Q\approx 0$. They found a positive polarization for $\Lambda$ but almost zero polarization for $\bar{\Lambda}$. On the other hand, the COMPASS collaboration COMPASS:2021bws found very small values consistent with zero for both $\Lambda$ and $\bar{\Lambda}$. The contribution from exclusive $s\bar{s}$ production, if any at all in these low energy experiments, will be significantly diluted by the inclusive background. On the other hand, diffractive events with a rapidity gap are copious at the EIC. By triggering on such events, one should be able to test our prediction in future.
[…]
Figure 3: Upper panels: Polarization of the strange (left) and antistrange (right) quarks in UPC at $W=30$ GeV, $Q=0$ GeV. Lower panels: Polarization of the charm (left) and anticharm (right) quarks in the same process. Here and in all the plots below, $k_{\perp}$ in the vertical axis is in units of GeV.
IV.2 Entanglement
The degree of entanglement between the quark and antiquark spins can be quantified by $\Delta_{1,2}$ introduced in (51). First, consider UPCs at the LHC and fix $W=500$ GeV, $Q=0$ and $k_{\perp}^{\rm max}=30$ GeV. In Fig. 4, we plot the left hand side of (50) in the $(z,k_{\perp})$ plane for $u\bar{u}$ (left), $c\bar{c}$ (middle) and $b\bar{b}$ (right) pairs, respectively. The results for $d\bar{d}$, $s\bar{s}$ pairs are similar to the $u\bar{u}$ case. In practice, these are plots of $\Delta_{2}^{T}$ that dominates in (50). Since $W$ is typically quite large at the LHC, the results are close to those in the Regge limit calculation in Fucilla:2025kit. In particular, the pairs are everywhere entangled (i.e., (50) is always satisfied). Light quarks are maximally entangled ($\Delta^{T}_{2}\approx 2$) along the line $z=\frac{1}{2}$ as discussed around (58). For massive quarks, an almost maximally entangled state
$|\Psi^{-}_{n}\rangle\approx\frac{1}{\sqrt{2}}\left(|+-\rangle_{n}-i|-+\rangle_{n}\right),$ (75)
is realized around $z=\frac{1}{2}$ and $k_{\perp}\approx m$. As explained in Fucilla:2025kit, this occurs because $I_{g}^{L}\approx 0$ (see the comment below (46)) and hence $C_{nn}^{T}\approx-1$ in this region. The peak is followed by a dip as $k_{\perp}$ increases (i.e., entanglement becomes weak) before rising again at higher $k_{\perp}$ where the mass effect becomes less important.
Next, we consider DIS at $W=30$ GeV and $Q=10$ GeV. In Fig. 5, we plot the left hand side of (50) (with $T\to{\rm DIS}$ in the superscript) for $u\bar{u}$ (left), $s\bar{s}$ (middle), $b\bar{b}$ (right) pairs. The $u\bar{u}$ plot is asymmetric under $z\leftrightarrow\bar{z}$. This is because the valence quark integrals $I_{L}^{q}$, $I_{T}^{q_{1,2}}$ with $q=u,d$ do not have this symmetry. Actually, the same is true in Fig. 4 (left), but the asymmetry is barely noticeable because at high energy the cross section is dominated by the gluon contribution which is symmetric. Since $Q$ is relatively high, the results reflect a significant interference between the transverse and longitudinal photon contributions. In $b\bar{b}$ (and $c\bar{c}$) production, we find narrow white regions where the pair is not entangled (i.e., the pair is separable). We have studied other sets of $(W,Q)$ and found that regions of no-entanglement appear only for $c\bar{c},b\bar{b}$ pairs in small patches at most. This means that, even though the conditions (53) do not hold away from the Regge limit, entanglement remains a robust feature of exclusive $q\bar{q}$ production.
[…]
Figure 4: Left hand side of (50) in UPC at $W=500$ GeV, $Q=0$. Left: $u\bar{u}$ pair, Middle: $c\bar{c}$ pair, Right: $b\bar{b}$ pair. The pairs are almost maximally entangled $\Delta_{2}^{T}\approx 2$ in the red region.
[…]
Figure 5: Left hand side (50) in DIS at $Q=10$ GeV, $W=30$ GeV. Left: $u\bar{u}$ pair, Middle: $s\bar{s}$ pair, Right: $b\bar{b}$ pair. The pair is not entangled in the white region.
IV.3 Bell nonlocality
Next, we delineate the regions where the $q\bar{q}$ pair exhibits Bell nonlocality. In Fig. 6 (left), we plot $\mu_{1}+\mu_{2}-1$ (see (60)) for $b\bar{b}$ pairs in the same UPC kinematics as in Fig. 4 (right). We see a large white region where $\mu_{1}+\mu_{2}-1<0$, meaning that the Bell-CHSH inequality is not violated there. This is in contrast to the finding in Fucilla:2025kit that, in the Regge limit, the pair exhibits Bell nonlocality in the entire kinematical region of $(z,k_{\perp})$. It is also in contrast to Fig. 4 (right) which shows that the pair remains entangled even away from the Regge limit. These observations are consistent with the fact that, in general, Bell nonlocality is a sufficient but not necessary condition for entanglement.
In the middle and right panels of Fig. 6, we plot the same quantity for the $s\bar{s}$ and $b\bar{b}$ pairs in DIS ($W=30$ GeV, $Q=10$ GeV) to be directly compared to the middle and right panels in Fig. 5. Again we find regions where Bell nonlocality is lost, even though the pairs are still entangled.
[…]
Figure 6: Plot of $\mu_{1}+\mu_{2}-1$, see (60). The Bell-CHSH inequality is not violated in the white regions. Left: $b\bar{b}$ production in UPC at $W=500$ GeV, $Q=0$, compare with Fig. 4 (right). Middle: $s\bar{s}$ production in DIS, compare with Fig. 5 (middle). Right: $b\bar{b}$ pair production in DIS, compare with Fig. 5 (right).
IV.4 Magic
Finally, in Fig. 7 and Fig. 8, we plot the stabilizer Rényi entropy $M_{2}$ (64) in UPC and DIS, respectively, for different quark flavors. As expected, $M_{2}$ is suppressed when entanglement is nearly maximal. But other than this general statement, it is difficult to anticipate in which kinematical regions $M_{2}$ is enhanced. We however see that $M_{2}$ tends to be larger for heavier quarks, with the maximal value around $0.58$ for the $b\bar{b}$ pair in DIS with $W=30$ GeV, $Q=3$ GeV, see Fig. 8 (right). Curiously, in the entire parameter space we have explored, $M_{2}$ is always smaller than (70), although in principle it can exceed this bound because $B_{n},\bar{B}_{n}\neq 0$.
[…]
Figure 7: Stabilizer Rényi entropy (64) in UPC ($W=500$ GeV) for $u\bar{u}$ (left), $c\bar{c}$ (middle), and $b\bar{b}$ (right) pairs.
[…]
Figure 8: Stabilizer Rényi entropy (64) in DIS ($W=30$ GeV, $Q=10$ GeV) for $s\bar{s}$ (left), $c\bar{c}$ (middle), and $b\bar{b}$ (right) pairs.
V Conclusions
In this paper, we have analyzed exclusive quark-antiquark production in DIS and UPCs from the viewpoint of Quantum Information Science (QIS). This is a generalization of the previous work in the high energy (Regge) limit Fucilla:2025kit to more practical CM energies at the EIC where collinear factorization with GPDs is an appropriate theoretical tool. The scattering amplitudes computed in the GPD framework are complex, in contrast to being purely real in the one-gluon exchange approximation Qi:2025onf and purely imaginary in the Regge limit. An immediate consequence is that massive $q\bar{q}$ pairs are transversely polarized without the usual $\alpha_{s}$ suppression in inclusive production. In low-energy UPC or DIS, the polarization can easily exceed 50% in certain kinematic regions. Moreover, we have observed rich patterns of entanglement, Bell nonlocality and magic due to the interference between the real and imaginary parts of the GPD amplitudes, and also between the transverse and longitudinal photons. Together with the EIC’s ability to vary $W$ and $Q$ over an unprecedentedly wide kinematical range, this makes DIS a particularly fascinating arena for exploring QIS in collider physics.
We have focused on the polarization and entanglement of the produced $q\bar{q}$ pairs. Whether these properties survive as observables in realistic experiments is a nontrivial problem that requires separate theoretical investigations. For massive quarks $s,c,b$ (or antiquarks $\bar{s},\bar{c},\bar{b}$), the canonical method is to tag events where the quarks fragment into heavy baryons $\Lambda,\Lambda_{c},\Lambda_{b}$ (or antibaryons $\bar{\Lambda},\bar{\Lambda}_{c},\bar{\Lambda}_{b}$) Galanti:2015pqa. Their subsequent weak decays such as $\Lambda\to p+\pi^{-}$ allows one to reconstruct the polarization vectors $B_{a},\bar{B}_{a}$ Bunce:1976yb; Lundberg:1989hw and the spin density matrix $C_{ab}$ Tornqvist:1980af; Afik:2025grr. For light quarks, methods outlined in Cheng:2025cuv may be applicable to DIS. In both cases, detailed simulations with a realistic detector setup at the EIC are necessary Lin:2025eci; Gu:2025ijz.
Acknowledgments
The authors were supported by the U.S. Department of Energy under Contract No. DE-SC0012704, and also by LDRD funds from Brookhaven Science Associates and the framework of the Saturated Glue (SURGE) Topical Theory Collaboration.
Appendix A Maximal magic without polarization
In this Appendix, we determine the upper limit of the stabilizer Rényi entropy Leone:2021rzd for generic two qubit systems with vanishing polarization. Following Liu:2025frx, we parametrize pure two-qubit states as
$|\psi\rangle=c_{1}|++\rangle+c_{2}|+-\rangle+c_{3}|-+\rangle+c_{4}|--\rangle,$ (76)
$c_{1}=\sin\theta_{1}\sin\theta_{2}e^{i\phi_{1}},\qquad c_{2}=\sin\theta_{1}\cos\theta_{2}e^{i\phi_{2}},\qquad c_{3}=\cos\theta_{1}\sin\theta_{3}e^{i\phi_{3}},\qquad c_{4}=\cos\theta_{1}\cos\theta_{3}.$ (77)
Without loss of generality, we may take $0\leq\theta_{1,2,3}\leq\frac{\pi}{2}$ and $0\leq\phi_{1,2,3}\leq 2\pi$. The polarization of the first qubit can be read off from the reduced density matrix traced over the second qubit
${\rm Tr}_{2}|\psi\rangle\langle\psi|=\begin{pmatrix}|c_{1}|^{2}+|c_{2}|^{2}&c_{1}c_{3}^{*}+c_{2}c_{4}^{*}\\ c_{1}^{*}c_{3}+c_{2}^{*}c_{4}&|c_{3}|^{2}+|c_{4}|^{2}\end{pmatrix}=\frac{1+\vec{B}_{1}\cdot\vec{\sigma}}{2}.$ (78)
If we require that $\vec{B}_{1}=0$, then
$|c_{1}|^{2}+|c_{2}|^{2}=|c_{3}|^{2}+|c_{4}|^{2}=\frac{1}{2},\qquad c_{1}c_{3}^{*}+c_{2}c_{4}^{*}=0.$ (79)
Similarly, if we require that $\vec{B}_{2}=0$,
$|c_{1}|^{2}+|c_{3}|^{2}=|c_{2}|^{2}+|c_{4}|^{2}=\frac{1}{2},\qquad c_{1}c_{2}^{*}+c_{3}c_{4}^{*}=0.$ (80)
Substituting (77) into (79), we get
$\sin\theta_{1}^{2}=\cos\theta_{1}^{2}\qquad\sin\theta_{2}\sin\theta_{3}e^{i(\phi_{1}-\phi_{3})}+\cos\theta_{2}\cos\theta_{3}e^{i\phi_{2}}=0.$ (81)
This means that
$\theta_{1}=\frac{\pi}{4}\qquad\phi_{1}-\phi_{3}=\phi_{2}\pm\pi,\qquad\theta_{3}=\frac{\pi}{2}-\theta_{2}.$ (82)
It is easy to check that when (82) is satisfied, (80) is automatically satisfied. The maximal value of $M_{2}$ for systems with $\vec{B}_{1}=\vec{B}_{2}=0$ is therefore obtained by minimizing (16) of Liu:2025frx with the constraints (82). There are many degenerate minima. For example, we find
$\cos\theta_{2}=\sqrt{\frac{1}{3}},\qquad\phi_{1}=\frac{\pi}{2},\qquad\phi_{2}=\frac{5}{4}\pi,\qquad\phi_{3}=\frac{\pi}{4},$ (83)
to a good numerical accuracy. The corresponding value of $M_{2}$ is given by (70).
Now let us further require that $\theta_{2}=\frac{\pi}{2}$, and accordingly, $\theta_{3}=0$. (Alternatively, one may set $\theta_{2}=0$ and $\theta_{3}=\frac{\pi}{2}$.) The density matrix becomes
$\rho^{L}=\frac{1}{2}\begin{pmatrix}1&0&0&e^{i\phi_{1}}\\ 0&0&0&0\\ 0&0&0&0\\ e^{-i\phi_{1}}&0&0&1\end{pmatrix}$. (84)
Minimizing (16) of Liu:2025frx under this additional constraint, we find the extremal condition $\cos(4\phi_{1})=-1$, or equivalently,
$\phi_{1}=\frac{\pi}{4},\frac{3\pi}{4},\frac{5\pi}{4},\frac{7\pi}{4}.$ (85)
At these points, $M_{2}$ takes the maximal value which coincides with (66). The four values of $\phi_{1}$ in (85) correspond to the four branches (68).